OpenStudy (anonymous):

How many real number solutions are there to the equation 0 = –3x² + x – 4?

OpenStudy (anonymous):

can you find the value of the discriminant?

OpenStudy (anonymous):

the value of what? i have no clue

OpenStudy (shayaan_mustafa):

when ever you required to know nature of roots then always always use discriminant. good luck . others are here to help you. good luck again.

OpenStudy (campbell_st):

for a quadratic \[ax^2 + bx + c = 0\] use the discriminant \[\Delta = b^2 - 4ac \] in your question a=-3, b = 1 and c= -4 then 1^2 - 4 x (-3) x (-4) = -47 since \[\Delta < 0\] no real solutions Rules \[\Delta > 0\] 2 unequal real solutions if \[\Delta\] is a square number the roots are rational. \[\Delta = 0\] repeated or equal roots and the senario above